OpenStudy (anonymous):

Solve for y. X=(4/5)y I need help!

OpenStudy (immanuelv):

what you need to do is divide both sides by (y) once you do that then you should divide that answer by x does that make sense?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

if you want to solve for y, you want to separate y and get it by itself. To do that, you have to move (4/5) Since your problem is x = (4/5)y It's basically x = (4/5) * y What is the opposite of multiplication? Division Divide (4/5) from y and x to get y by iself When you divide a fraction, it's cross multiply x / (4/5) is basically x * (5/4) and (4/5) cancel out with (4/5)y so now y is alone, rewrite your equation it should be (5/4)x = y